# Thinking Outside the (Gray) Box: A Context-Based Score for Assessing Value and Originality in Neural Text Generation

### v2

@@ A Context-Based Score for Valuable and Original Generation
We now discuss the implementation of the CoVO score with autoregressive models. Using the notation introduced above, in the context of a $\boldsymbol{\theta}$-parameterized LLM, $p(\mathbf{y}|\mathbf{x})$ can be expressed as $\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})$. However, considering just the product of all the conditioned probabilities for an optimization problem would lead to preferring shorter sequences. To avoid this, we propose to use the $T$-th root: $T]{\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})}$. By leveraging the properties of the logarithm, we obtain:
$s_{CoVO}^{AR}$ $=\lambda_{v}\frac{\sum_{i=1}^{|\mathbf{x}|}\log p_{\boldsymbol{\theta}}(x_{i}|\mathbf{x}_{<i},\mathbf{y})}{|\mathbf{x}|}$ (10)
$-\lambda_{o}\frac{\sum_{j=1}^{|\mathbf{y}|}\log p_{\boldsymbol{\theta}}(y_{j}|\mathbf{y}_{<j},\mathbf{x})}{|\mathbf{y}|}.$
@@
It is worth noting that the vocabulary of an LLM can be extremely large, which can cause $p_{\boldsymbol{\theta}}(a|b)$ to be small even when $a$ is the most probable event given $b$. In particular, when an LLM generates $\mathbf{y}$ given $\mathbf{x}$ and then evaluates both $p_{\boldsymbol{\theta}}(\mathbf{y}|\mathbf{x})$ and $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$, this can lead to a significant discrepancy between the magnitude of value and diversity. Since $\mathbf{y}$ has been sampled from a model based on $p_{\boldsymbol{\theta}}$, its probability would be high by definition. However, there may be various ways (possibly through synonyms) to define $\mathbf{y}$, leading to a smaller probability of $\mathbf{x}$.
@@
Inspired by Macedo, Reisenzein, and Cardoso (2004), we propose to counteract this problem by normalizing $p_{\boldsymbol{\theta}}(a|b)$ via $n^{\prime}=\frac{n-n_{min}}{n_{max}-n_{min}}$. For probabilities, $n_{min}=0$, while $n_{max}=\max p_{\boldsymbol{\theta}}(b)$, thus obtaining the overall mapping for $p_{\boldsymbol{\theta}}$: $\frac{p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})}{\max p_{\boldsymbol{\theta}}(\mathbf{y}_{<t},\mathbf{x})}$. Once again, by applying the properties of logarithms, we obtain:
$s_{CoVO}^{AR_{norm}}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})=\lambda_{v}s_{v}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})+\lambda_{o}s_{o}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})=$ (11)
$\lambda_{v}\frac{\sum_{i=1}^{|\mathbf{x}|}(\log p_{\boldsymbol{\theta}}(x_{i}|\mathbf{x}_{<i},\mathbf{y})-\max\log p_{\boldsymbol{\theta}}(\mathbf{x}_{<i},\mathbf{y}))}{|\mathbf{x}|}-$
$\lambda_{o}\frac{\sum_{j=1}^{|\mathbf{y}|}(\log p_{\boldsymbol{\theta}}(y_{j}|\mathbf{y}_{<j},\mathbf{x})-\max\log p_{\boldsymbol{\theta}}(\mathbf{y}_{<j},\mathbf{x}))}{|\mathbf{y}|}.$
@@
Calculating $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$ is not trivial. Since LLMs are trained to complete text sequences, it is unlikely that they would generate the source text immediately after the target text (which, we should remember, is generated immediately after the source text). To address this, we consider an approximation $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y}^{\prime})$, where $\mathbf{y}^{\prime}=\mathbf{y}+\mathbf{q}$. Here, $\mathbf{q}$ represents an additional question, such as “How would you describe this text?” or a similar formulation designed solely to increase the likelihood of generating the source text $\mathbf{x}$ (as well as alternative sources). Once the CoVO score has been defined, its adoption in an RL framework is straightforward. As previously introduced, we can directly utilize our CoVO score as the final reward for the generated sequence.
@@
Then, the model can be trained with any policy gradient method. Our experiments leverage GRPO (Shao et al. 2024), which is a state-of-the-art choice for training language models. As introduced above, GRPO adds a per-token KL divergence term to the loss rather than to the single rewards. Usually, the KL divergence is approximated with the following estimator (Schulman 2020): $r_{ref}(\boldsymbol{\theta})-\log r_{ref}(\boldsymbol{\theta})-1$, where $r_{ref}(\boldsymbol{\theta})=\frac{\pi_{ref}}{\pi_{\boldsymbol{\theta}}}$ and $\pi_{ref}$ is the reference policy, i.e., the model before GRPO training. In particular, GRPO aims to minimize the KL divergence, thus the second term $-\log r_{ref}=-\log\pi_{ref}+\log\pi_{\boldsymbol{\theta}}$ can be seen as made of two optimization problems: the minimization of the originality component $-\log\pi_{ref}$ from Equation 9, and the maximization of surprisal, or self-information, $-\log\pi_{\boldsymbol{\theta}}$ under the current model. In other words, this second term somehow trades off a portion of the originality component under the reference model (proportional to the $\beta\ll 1$ coefficient) with the surprisal under the current model. However, the first term of the KL approximation keeps the two policies closer, preventing the trained one from deviating too much and potentially disrupting the impact of the originality component in favor of the value component.
Experiments
@@
. In all experiments, we employ two settings, i.e., GRPO to maximize the score from Equation 11 without the KL divergence loss, i.e., with $\beta=0.0$ (CoVO); and GRPO to maximize the score from Equation 11 with the KL divergence loss, i.e., with $\beta=0.05$ (CoVO + KL). Both methods assume $\lambda_{v}=\lambda_{o}=1.0$.
Poetry Generation
@@
Meta-Llama-3-8B 0.9870.9870.987 0.3000.3000.300 / 0.1770.1770.177 6.0676.0676.067 / 191919 $0.664_{\pm 0.069}$ 1.0001.0001.000 0.4440.4440.444 / 0.1320.1320.132 5.8535.8535.853 / 686868 $0.621_{\pm 0.063}$
+ CoVO 1.0001.0001.000 0.2000.2000.200 / 0.0840.0840.084 4.8804.8804.880 / 0777 $0.749_{\pm 0.055}$ 0.9600.9600.960 0.2670.2670.267 / 0.0460.0460.046 4.9334.9334.933 / 0666 $0.661_{\pm 0.060}$
+ CoVO + KL 0.9870.9870.987 0.3330.3330.333 / 0.0660.0660.066 5.1075.1075.107 / 0999 $0.745_{\pm 0.057}$ 0.9730.9730.973 0.2890.2890.289 / 0.1170.1170.117 5.1605.1605.160 / 222222 $0.682_{\pm 0.059}$
Table 1: Aggregate results of CoVO scores at inference time considering both training prompts (left) and testing prompts (right). Scores on the poetical metrics are reported at the line level (L) and syllable level (S) and only consider requests for styles with specific metrical properties. Under T-LCS, we report both the mean and the maximum longest common substring across all generated poems. The mean and the 95% confidence interval are reported for tone adherence.
Experimental Setup.
[…]
Experimental Results.
Table 1 reports the scores about the compliance of poetical constraints at the syllable and line levels, lexical correctness (as the ratio of poems not containing noisy text), tone adherence (as the zero-shot classification of that poem being of that tone rather than its opposite), and accidental reproduction rate (as the mean and maximum token-based longest common substring).
@@
Interestingly, these considerations align well with our CoVO score. Figure 2 reports the value and originality according to Equation 11 under the pre-trained model. While the two methods do not significantly differ from the baseline (which is possibly due to the opposite forces of value and originality (Varshney 2019)), we again see that the presence of KL leads to slightly higher value, while its absence leads to slightly higher originality. However, aggregated scores, such as those presented here, might be insufficient. For a more complete overview, we also conducted a fine-grained analysis of the generated poems in the Technical Appendix.
Figure 2: The distribution of value and originality (according to our scores) for the in-distribution and out-of-distribution poems generated by the baseline and our four methods.
Math Problem Resolution
Experimental Setup.
@@
The second set of experiments concerns a more practical and quantitative task, as it aims to teach the LLM to solve mathematical problems through more diverse procedures. In particular, we focus on the Mistral-based (Jiang et al. 2023) MetaMath-Mistral-7B model, i.e., fine-tuned with self-supervised learning on the MetaMathQA (Yu et al. 2024). It is a dataset of textual math questions paired with responses where the numerical answer is easily separable from the textual procedure. While the entire set contains 395k entries, making an additional training epoch too expensive, MetaMathQA is composed of entries from two different training sets, then augmented with various techniques: GSM8K (Cobbe et al. 2021) and MATH (Hendrycks et al. 2021). Since we are only interested in the questions, we limit our training to a single epoch over those datasets. Moreover, we exclude all questions with a tokenized length of either question or answer greater than 512, obtaining 14876 out of 14973 total entries.
We separate the procedure and the answer from each solution to train our model and use the numerical answer to check the correctness of the predicted solution. Because of this, the RL problem can be formulated considering up to two rewards: our CoVO score computed on the procedure and an extrinsic reward based on the correctness of the answer. As for the previous case study, instead of fine-tuning the entire model, we adopt a more parameter-efficient strategy with LoRA and use the original model to perform the CoVO score computation. Following (Yu et al. 2024), the outputs are obtained with a greedy strategy.
@@
The evaluation considers both GSM8K and MATH test sets (limited to the entries with a tokenized length of question and answer smaller than 512, i.e., all 1319 entries for GSM8K and 4546 out of 5000 for MATH). We compute the percentage of correct solutions together with two diversity metrics: expectation-adjusted distinct N-grams (EAD) (Liu et al. 2022) and sentence embedding cosine similarity (SBERT) (Hong et al. 2024), which should measure syntactical and semantical diversity, respectively (Kirk et al. 2024). EAD counts the number of distinct N-grams (averaging over $N=1\ldots 5$) across all generated responses and removes the bias toward shorter inputs by scaling the number of distinct tokens based on their expectations. The SBERT metric computes the average of the cosine similarity between the embeddings of any possible pairs of outputs and returns 1 minus the similarity. This was originally based on Sentence-BERT (Reimers and Gurevych 2019), we employ instead the more recent all-mpnet-base-v2, as suggested by their developers [^2] https://huggingface.co/sentence-transformers/bert-large-nli-stsb-mean-tokens.
@@
[…]
+ CoVO + Ext $\mathbf{78.33}\%(4)$ 2.03402.03402.0340 $\mathbf{0.6404}$ $0.0628_{\pm.0056}$ $0.0088_{\pm.0011}$
+ CoVO + Ext + KL $\underline{77.95}\%(4)$ 2.03672.03672.0367 0.64020.64020.6402 $0.0638_{\pm.0057}$ $0.0089_{\pm.0011}$
@@
[…]
+ CoVO + Ext $\mathbf{33.76}\%(503)$ 5.81145.81145.8114 0.80300.80300.8030 $0.1136_{\pm.0047}$ $0.0259_{\pm.0015}$
[…]
@@
Experimental Results.
Tables 2 and 3 report the results for the GSM8K and MATH test sets. For the GSM8K test set, while all methods achieve similar results, using the CoVO score only (with and without the KL loss) leads to greater EAD diversity and to diverge more from the original model, while the presence of the math reward leads to greater accuracy, especially without KL.
@@
NoveltyBench
Experimental Setup.
Finally, we also experiment with NoveltyBench (Zhang et al. 2025), a very recent benchmark that aims to evaluate the ability of language models to produce multiple distinct and high-quality outputs. NoveltyBench contains two sets of prompts thought for eliciting diverse responses: the ‘curated’ partition, with 100 prompts manually curated by the paper’s authors, and the ‘wildchat’ partition, with 1000 prompts sourced from the WildChat-1M dataset (Zhao et al. 2024). The benchmark requires a language model to generate 10 outputs per prompt using a temperature setting of 1.01.01.0, after which it computes scores for novelty and quality. To assess novelty, the 10 outputs are grouped into equivalence classes using a fine-tuned DeBERTa model (He et al. 2021), and the number of distinct classes is reported as the novelty score. For the latter, it computes the cumulative utility of the 10 outputs, where the utility is zero if the $i$-th output has the same equivalence class as a precedent output, and a calibrated reward from Skywork-Reward-Gemma-2-27B-v0.2 (Liu et al. 2024) otherwise (we refer to Zhang et al. (2025) for full details). To evaluate our approach with NoveltyBench, we fine-tune the Llama-3.2-3B-Instruct model (Grattafiori et al. 2024) for a single epoch on the ‘wildchat’ partition (the full parameters are reported in the Technical Appendix); then, we compute the novelty and quality scores on the ‘curated’ partition and we compare them with the performance of the original model.
Experimental Results.
Figure 3 reports the quality and novelty scores achieved by our methods, compared to those from the original instructed model. Optimizing for the CoVO score results in substantial improvements in both novelty and quality metrics, with the greatest gains in novelty. Moreover, these results underscore the interplay between KL loss and our CoVO score: incorporating the KL penalty tends to improve quality, but at the cost of reduced novelty.
Discussion

### v3

@@ 5 Implementation and Optimization with Autoregressive Models
We now discuss the implementation of the CoVO score with autoregressive models. Using the notation introduced above, in the context of a $\bm{\theta}$-parameterized LLM, $p(\mathbf{y}|\mathbf{x})$ can be expressed as $\prod_{t=1}^{T}p_{\bm{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})$. However, considering just the product of all the conditioned probabilities for an optimization problem would lead to preferring shorter sequences. To avoid this, we propose to use the $T$-th root: $\sqrt[T]{\prod_{t=1}^{T}p_{\bm{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})}$. By leveraging the properties of the logarithm, we obtain:
$s_{CoVO}^{AR}=\lambda_{v}\frac{\sum_{i=1}^{|\mathbf{x}|}\log p_{\bm{\theta}}(x_{i}|\mathbf{x}_{<i},\mathbf{y})}{|\mathbf{x}|}-\lambda_{o}\frac{\sum_{j=1}^{|\mathbf{y}|}\log p_{\bm{\theta}}(y_{j}|\mathbf{y}_{<j},\mathbf{x})}{|\mathbf{y}|}.$ (10)
@@
It is worth noting that the vocabulary of an LLM can be extremely large, which can cause $p_{\bm{\theta}}(a|b)$ to be small even when $a$ is the most probable event given $b$. In particular, when an LLM generates $\mathbf{y}$ given $\mathbf{x}$ and then evaluates both $p_{\bm{\theta}}(\mathbf{y}|\mathbf{x})$ and $p_{\bm{\theta}}(\mathbf{x}|\mathbf{y})$, this can lead to a significant discrepancy between the magnitude of value and diversity. Since $\mathbf{y}$ has been sampled from a model based on $p_{\bm{\theta}}$, its probability would be high by definition. However, there may be various ways (possibly through synonyms) to define $\mathbf{y}$, leading to a smaller probability of $\mathbf{x}$.
@@
Inspired by Macedo et al. (2004), we propose to counteract this problem by normalizing $p_{\bm{\theta}}(a|b)$ via $n^{\prime}=\frac{n-n_{min}}{n_{max}-n_{min}}$. For probabilities, $n_{min}=0$, while $n_{max}=\max_{v\in\mathcal{V}}p_{\bm{\theta}}(b)$, thus obtaining the overall mapping for $p_{\bm{\theta}}$: $\frac{p_{\bm{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})}{\max_{v\in\mathcal{V}}p_{\bm{\theta}}(\mathbf{y}_{<t},\mathbf{x})}$. Once again, by applying the properties of logarithms, we obtain:
$s_{CoVO}^{AR_{norm}}(\mathbf{x},\mathbf{y},p_{\bm{\theta}})=\lambda_{v}s_{v}(\mathbf{x},\mathbf{y},p_{\bm{\theta}})+\lambda_{o}s_{o}(\mathbf{x},\mathbf{y},p_{\bm{\theta}})=$ (11)
$\lambda_{v}\frac{\sum_{i=1}^{|\mathbf{x}|}(\log p_{\bm{\theta}}(x_{i}|\mathbf{x}_{<i},\mathbf{y})-\max_{v\in\mathcal{V}}\log p_{\bm{\theta}}(\mathbf{x}_{<i},\mathbf{y}))}{|\mathbf{x}|}$-
$\lambda_{o}\frac{\sum_{j=1}^{|\mathbf{y}|}(\log p_{\bm{\theta}}(y_{j}|\mathbf{y}_{<j},\mathbf{x})-\max_{v\in\mathcal{V}}\log p_{\bm{\theta}}(\mathbf{y}_{<j},\mathbf{x}))}{|\mathbf{y}|}.$
@@
Calculating $p_{\bm{\theta}}(\mathbf{x}|\mathbf{y})$ is not trivial. Since LLMs are trained to complete text sequences, it is unlikely that they would generate the source text immediately after the target text (which, we should remember, is generated immediately after the source text). To address this, we consider an approximation $p_{\bm{\theta}}(\mathbf{x}|\mathbf{y}^{\prime})$, where $\mathbf{y}^{\prime}=\mathbf{y}+\mathbf{q}$. Here, $\mathbf{q}$ represents an additional question, such as “How would you describe this text?” or a similar formulation designed solely to increase the likelihood of generating the source text $\mathbf{x}$ (as well as alternative sources).
@@
Once the CoVO score has been defined, its adoption in an RL framework is straightforward. As previously introduced, we can directly utilize our CoVO score as the final reward for the generated sequence. Then, the model can be trained with any policy gradient method. Our experiments leverage GRPO (Shao et al., 2024), which is a state-of-the-art choice for training language models. As introduced above, GRPO adds a per-token KL divergence term to the loss rather than to the single rewards. Usually, the KL divergence is approximated with the following estimator (Schulman, 2020): $r_{ref}(\bm{\theta})-\log r_{ref}(\bm{\theta})-1$, where $r_{ref}(\bm{\theta})=\frac{\pi_{ref}}{\pi_{\bm{\theta}}}$ and $\pi_{ref}$ is the reference policy, i.e., the model before GRPO training. In particular, GRPO aims to minimize the KL divergence, thus the second term $\log r_{ref}=-\log\pi_{ref}+\log\pi_{\bm{\theta}}$ can be seen as made of two optimization problems: the minimization of the originality component $\log\pi_{ref}$ from Equation 9, and the maximization of surprisal, or self-information, $\log\pi_{\bm{\theta}}$ under the current model. In other words, this second term somehow trades off a portion of the originality component under the reference model (proportional to the $\beta\ll 1$ coefficient) with the surprisal under the current model. However, the first term of the KL approximation keeps the two policies closer, preventing the trained one from deviating too much and potentially disrupting the impact of the originality component in favor of the value component.
6 Experiments
@@
. In all experiments, we employ two settings, i.e., GRPO to maximize the score from Equation 11 without the KL divergence loss, i.e., with $\beta=0.0$ (CoVO); and GRPO to maximize the score from Equation 11 with the KL divergence loss, i.e., with $\beta=0.05$ (CoVO + KL). Both methods assume $\lambda_{v}=\lambda_{o}=1.0$. While it is common to induce diversity at the sampling level (e.g., through min-$p$ (Minh et al., 2025) or conformative decoding (Peeperkorn et al., 2025)), we restrict our baselines to the original model and, if available, a model tuned solely on the environmental reward. Our approach is orthogonal to the chosen sampling strategy and can benefit from more advanced methods. In our evaluation, we aim at verifying whether our reward scheme can increase value and originality, regardless of how the output is sampled.
6.1 Poetry Generation
@@
Meta-Llama-3-8B $0.987$ $0.300$ / $0.177$ $6.067$ / 1919 $0.664_{\pm 0.069}$ $1.000$ $0.444$ / $0.132$ $5.853$ / $68$ $0.621_{\pm 0.063}$
+ CoVO $1.000$ $0.200$ / $0.084$ $4.880$ / 077 $0.749_{\pm 0.055}$ $0.960$ $0.267$ / $0.046$ $4.933$ / 066 $0.661_{\pm 0.060}$
+ CoVO + KL $0.987$ $0.333$ / $0.066$ $5.107$ / 099 $0.745_{\pm 0.057}$ $0.973$ $0.289$ / $0.117$ $5.160$ / $22$ $0.682_{\pm 0.059}$
Table 1: Aggregate results of generated poems considering both training prompts (left) and testing prompts (right). Scores on the poetical metrics are reported at the line level (L) and syllable level (S) and only consider requests for styles with specific metrical properties. Under T-LCS, we report both the mean and the maximum longest common substring across all generated poems. The mean and the 95% confidence interval are reported for tone adherence.
Figure 2: The distribution of value and originality (according to our scores) for the in-distribution and out-of-distribution poems generated by the baseline and our two methods.
Experimental Setup. The first set of experiments concerns a very common creative task, aiming to teach the LLM to generate poems that are both more original and valuable. More specifically, we follow the approach outlined by Bradley et al. (2024) and instruct the model to write a poem in a particular style and tone. We consider the Meta-Llama-3-8B model (Grattafiori et al., 2024) as our pre-trained agent. Since we do not use the instruction-tuned model, we prompt it with some few-shot examples of the task to make it more likely to produce the desired output in the desired form (the full prompt is reported in Appendix A, together with the full training parameters). Instead of fine-tuning the entire network, we consider Low-Rank Adaptation (LoRA) (Hu et al., 2022). The original model is also used to compute the score. All sampling happens with top-$k$ ($k=50$) and at a temperature of $1.0$. We perform a quantitative evaluation where we compute poetical metrics for quality (lexical correctness of poems, adherence to line- and syllable-level constraints, and tone adherence to the request through zero-shot classification (Yin et al., 2019) with bart-large-mnli model (Lewis et al., 2020)) and for originality (accidental reproduction of existing poems). For the latter, we define a Token-based Longest Common Substring (T-LCS) score, and we use it by comparing generated poems with a reference dataset of approx. 84k public-domain poems extracted from Project Gutenberg (please refer to Appendix B for a first presentation of our GutenVerse dataset). While a generated poem can be an accidental reproduction of a protected work or a different kind of text (e.g., a song), we believe it can provide a useful evaluation tool to understand the general degree of originality.
Experimental Results. Table 1 reports the scores about the compliance of poetical constraints at the syllable and line levels, lexical correctness (as the ratio of poems not containing noisy text), tone adherence (as the zero-shot classification of that poem being of that tone rather than its opposite), and accidental reproduction rate (as the mean and maximum token-based longest common substring).
@@
Interestingly, these considerations align well with our CoVO score. Figure 2 reports the value and originality according to Equation 11 under the pre-trained model. While the two methods do not significantly differ from the baseline (which is possibly due to the opposite forces of value and originality (Varshney, 2019)), we again see that the presence of KL leads to slightly higher value, while its absence leads to slightly higher originality. However, aggregated scores, such as those presented here, might be insufficient. For a more complete overview, we also conducted a fine-grained analysis of the generated poems in Appendix C.
6.2 Math Problem Resolution
@@
Experimental Setup. The second set of experiments concerns a more practical and quantitative task, as it aims to teach the LLM to solve mathematical problems through more diverse procedures. In particular, we focus on the Mistral-based (Jiang et al., 2023) MetaMath-Mistral-7B model, i.e., fine-tuned with self-supervised learning on the MetaMathQA (Yu et al., 2024). It is a dataset of textual math questions paired with responses where the numerical answer is easily separable from the textual procedure. While the entire set contains 395k entries, making an additional training epoch too expensive, MetaMathQA is composed of entries from two different training sets, then augmented with various techniques: GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). Since we are only interested in the questions, we limit our training to a single epoch over those datasets. Moreover, we exclude all questions with a tokenized length of either question or answer greater than 512, obtaining 14876 out of 14973 total entries. We separate the procedure and the answer from each solution to train our model and use the numerical answer to check the correctness of the predicted solution. The RL problem can then be formulated considering up to two rewards: our CoVO score computed on the procedure and a verifiable, extrinsic reward based on the correctness of the answer. Instead of fine-tuning the entire model, we adopt a more parameter-efficient strategy with LoRA and use the original model to perform the CoVO score computation. Following Yu et al. (2024), the outputs are obtained with a greedy strategy.
@@
The evaluation considers both GSM8K and MATH test sets (limited to the entries with a tokenized length of question and answer smaller than 512, i.e., all 1319 entries for GSM8K and 4546 out of 5000 for MATH). We compute the percentage of correct solutions together with two diversity metrics: expectation-adjusted distinct N-grams (EAD) (Liu et al., 2022) and sentence embedding cosine similarity (SBERT) (Hong et al., 2024), which should measure syntactical and semantical diversity, respectively (Kirk et al., 2024). EAD counts the number of distinct N-grams (averaging over $N=1\ldots 5$) across all generated responses and removes the bias toward shorter inputs by scaling the number of distinct tokens based on their expectations. The SBERT metric computes the average of the cosine similarity between the embeddings of any possible pairs of outputs and returns 1 minus the similarity. This was originally based on Sentence-BERT (Reimers & Gurevych, 2019), we employ instead the more recent all-mpnet-base-v2, as suggested by their developers [^2] https://huggingface.co/sentence-transformers/bert-large-nli-stsb-mean-tokens.
@@
MetaMath-Mistral-7B $77.96\%(3)$ $2.0071$ $0.6402$ - -
+ Ext $78.18\%(4)$ $\underline{2.0045}$ $\mathbf{0.6404}$ $\underline{0.0081_{\pm.0021}}$ $\underline{0.0008_{\pm.0002}}$
+ Ext + KL $78.08\%(5)$ $2.0077$ $\underline{0.6401}$ $\underline{0.0096_{\pm.0024}}$ $\underline{0.0011_{\pm.0004}}$
+ CoVO $78.12\%(3)$ $\mathbf{2.0509}$ $0.6403$ $\mathbf{0.0854_{\pm.0062}}$ $\mathbf{0.0118_{\pm.0012}}$
+ CoVO + KL $78.12\%(3)$ $2.0464$ $0.6402$ $\mathbf{0.0879_{\pm.0063}}$ $\mathbf{0.0122_{\pm.0013}}$
+ CoVO + Ext $\mathbf{78.33}\%(4)$ $2.0340$ $\mathbf{0.6404}$ $0628_{\pm.0056} 0.0$$0088_{\pm.0011}$
+ CoVO + Ext + KL $\underline{77.95}\%(4)$ $2.0367$ $0.6402$ $0638_{\pm.0057} 0.0$$0089_{\pm.0011}$
@@
+ Ext + KL $33.56\%(476)$ $5.7517$ $0.8028$ $\underline{0.0345_{\pm.0030}}$ $\underline{0.0075_{\pm.0008}}$
+ CoVO $33.29\%(521)$ $5.8219$ $0.8029$ $\mathbf{0.1457_{\pm.0049}}$ $\mathbf{0.0339_{\pm.0016}}$
+ CoVO + KL $32.79\%(533)$ $\mathbf{5.8442}$ $0.8030$ $\mathbf{0.1479_{\pm.0050}}$ $\mathbf{0.0343_{\pm.0016}}$
+ CoVO + Ext $\mathbf{33.76}\%(503)$ $5.8114$ $0.8030$ $1136_{\pm.0047} 0.0$$0259_{\pm.0015}$
+ CoVO + Ext + KL $33.74\%(497)$ $5.8082$ $0.8031$ $1102_{\pm.0047} 0.0$$0250_{\pm.0015}$
@@
Experimental Results. Tables 2 and 3 report the results for the GSM8K and MATH test sets. For the GSM8K test set, while all methods achieve similar results, using the CoVO score only (with and without the KL loss) leads to greater EAD diversity and to diverge more from the original model, while the presence of the math reward leads to greater accuracy, especially without KL.
@@
6.3 NoveltyBench
Experimental Setup. Finally, we also experiment with NoveltyBench (Zhang et al., 2025), a very recent benchmark that aims to evaluate the ability of language models to produce multiple distinct and high-quality outputs. NoveltyBench contains two sets of prompts thought for eliciting diverse responses: the ‘curated’ partition, with 100 prompts manually curated by the paper’s authors, and the ‘wildchat’ partition, with 1000 prompts sourced from the WildChat-1M dataset (Zhao et al., 2024). The benchmark requires a language model to generate 10 outputs per prompt using a temperature setting of $1.0$, after which it computes scores for novelty and quality. To assess novelty, the 10 outputs are grouped into equivalence classes using a fine-tuned DeBERTa model (He et al., 2021), and the number of distinct classes is reported as the novelty score. To assess quality, it computes the cumulative utility of the 10 outputs, where the utility is zero if the $i$-th output has the same equivalence class as a precedent output, and a calibrated reward from Skywork-Reward-Gemma-2-27B-v0.2 (Liu et al., 2024) otherwise (we refer to Zhang et al. (2025) for full details). To evaluate our approach with NoveltyBench, we fine-tune the Llama-3.2-3B-Instruct model (Grattafiori et al., 2024) for a single epoch on the ‘wildchat’ partition (the full parameters are reported in Appendix A); then, we compute the novelty and quality scores on the ‘curated’ partition.
Experimental Results. Figure 3 reports the quality and novelty scores achieved by our methods, compared to those from the original instructed model. Optimizing for the CoVO score results in substantial improvements in both novelty and quality metrics, with the greatest gains in novelty. Moreover, these results underscore the interplay between KL loss and our CoVO score: incorporating the KL penalty tends to improve quality, but at the cost of reduced novelty.
7 Discussion